OpenStudy (anonymous):

Solve for x: |x + 2| + 16 = 14 x = −32 and x = −4 x = −4 and x = 0 x = 0 and x = 28 No solution Will fan and medal!

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ |x + 2| + 16 = 14 }\) subtract 16 from both sides, \(\normalsize\color{blue}{ |x + 2| + 16 \color{red}{ -16 }= 14\color{red}{ -16 } }\) \(\normalsize\color{blue}{ |x + 2| = -2 }\) this gives, \(\normalsize\color{blue}{ x + 2 = -2 ~~~~~~or~~~~~~ x+2=2 }\)

OpenStudy (solomonzelman):

can you finish the problem ?

OpenStudy (anonymous):

How?

OpenStudy (solomonzelman):

solve each equation

OpenStudy (anonymous):

Idk how to..

OpenStudy (anonymous):

Can you explain how?

OpenStudy (solomonzelman):

can you first solve \(\normalsize\color{blue}{ x + 2 = -2 }\) and then solve \(\normalsize\color{blue}{ x + 2 = 2 }\) ?

OpenStudy (solomonzelman):

in each equation, you subtract 2 from both sides.

OpenStudy (anonymous):

\[2-2 & 2-(-2)\] ?

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ x + 2 = -2~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~or~~~~~~~~~~~~~~~~~x+2=2 }\) \(\normalsize\color{blue}{ x + 2 \color{red}{ -2 }= -2\color{red}{ -2 }~~~~~~~~~~~~~~~~~or~~~~~~~~~~~~~~~~~x+2\color{red}{ -2 }=2\color{red}{ -2 } }\) \(\normalsize\color{blue}{ x = ?~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~or~~~~~~~~~~~~~~~~~x=? }\)